OpenStudy (anonymous):

What are the possible number of positive, negative, and complex zeros of f(x) = 2x3 – 5x2 – 6x + 4 ?

OpenStudy (ranga):

Apply Descartes' Rule of Signs.

OpenStudy (anonymous):

i Know that, but I'm confused on whether I got it right or not? @ranga

OpenStudy (ranga):

And what did you get?

OpenStudy (anonymous):

I got 1 positive 1 negative and I'm not sure about the complex zeros

OpenStudy (ranga):

f(x) = 2x^3 – 5x^2 – 6x + 4 Coefficients are: +2 -5 -6 +4 How many times does the coefficient change signs?

OpenStudy (anonymous):

twice

OpenStudy (anonymous):

That would be 2 positives?

OpenStudy (ranga):

It changes twice is correct. That means the number of possible positive roots are: 2 or 0. (you have to decrease by 2 each time until it reaches 0 or 1. If it changes sign 6 times then possible positive roots are: 6, 4, 2, 0.)

OpenStudy (anonymous):

Ok that's understanding, so negative would be 1 or 0?

OpenStudy (ranga):

For negative roots, first replace x by -x in f(x): f(x) = 2x^3 – 5x^2 – 6x + 4 f(-x) = 2(-x)^3 - 5(-x)^2 - 6(-x) + 4 f(-x) = -2x^3 - 5x^2 + 6x + 4 Coefficients: -2 -5 +6 +4 How many times does the coefficient change signs?

OpenStudy (anonymous):

Once , 1 or 0 negative roots?

OpenStudy (ranga):

It changes sign once is correct. The number of possible negative roots is 1. (you cannot decrease the number of roots by 2 anymore because you have to stop when it reaches 0 or 1. here it is already 1 and so it stops.) So there is no ambiguity here. This function definitely has ONE negative root. But with positive it can have 2 or 0 roots.

OpenStudy (anonymous):

wait when you substituted x by -x doesn't two negatives change to a positive?

OpenStudy (ranga):

f(x) = 2x^3 – 5x^2 – 6x + 4 f(-x) = 2(-x)^3 - 5(-x)^2 - 6(-x) + 4 f(-x) = -2x^3 - 5x^2 + 6x + 4 Coefficients: -2 -5 +6 +4 Which step you have a question?

OpenStudy (anonymous):

For the 2nd and 3rd, if it's -5(-x)^2 wouldn't the coefficient be a positive 5

OpenStudy (ranga):

The -5 is not squared. Only the x is replaced by -x. So x^2 becomes (-x)^2 = +x^2 -5(x^2) replace x by -x -5(-x)^2 = -5 * x^2 = -5x^2

OpenStudy (anonymous):

so it makes no difference that's why 5 is still a negative because -6(-x) became a positive 6

OpenStudy (ranga):

f(x) = 2x^3 – 5x^2 – 6x + 4 Do it term by term. First term is 2x^3. Replace x by -x: 2(-x)^3 = -2x^3 Second term: -5x^2. Replace x by -x: -5(-x)^2 = -5x^2 Third term: -6x. Replace x by -x: -6(-x) = +6x Fourth term: +4

OpenStudy (anonymous):

Okay I got it. Every other term? and what about complex zeros?

OpenStudy (ranga):

When you replace x by -x, only the ODD powered x will change the sign of the coefficient. The EVEN power will retain the same sign.

OpenStudy (ranga):

This is a third degree polynomial and so it should have a total of 3 roots. We already know that it definitely has one negative root. Positive root can be 2 or 0. But in any case the total number of roots should always add up to 3. List the possibilities: Negative Positive Complex Total 1 2 - 3 1 0 2 3 Those are the two possibilities: a) 1 negative, 2 positive or b) 1 negative, 2 complex.

OpenStudy (anonymous):

Okay so 2 or 0 complex. correct?

OpenStudy (ranga):

correct. 1 negative, 2 or 0 positive, 0 or 2 complex.

OpenStudy (anonymous):

Thank you so much ( :

OpenStudy (ranga):

You are welcome.